# Common Corpus: The Largest Collection of Ethical Data for LLM Pre-Training

### v1

@@ Abstract
Large Language Models (LLMs) are pre-trained on large amounts of data from different sources and domains. These data most often contain trillions of tokens with large portions of copyrighted or proprietary content, which hinders the usage of such models under AI legislation. This raises the need for truly open pre-training data that is compliant with the data security regulations. In this paper, we introduce Common Corpus [^1] https://huggingface.co/datasets/PleIAs/common_corpus, the largest open dataset for language model pre-training. The data assembled in Common Corpus are either uncopyrighted or under permissible licenses and amount to about two trillion tokens. The dataset contains a wide variety of languages, ranging from the main European languages to low-resource ones rarely present in pre-training datasets; in addition, it includes a large portion of code data. The diversity of data sources in terms of covered domains and time periods opens up the paths for both research and entrepreneurial needs in diverse areas of knowledge. In this technical report, we present the detailed provenance of data assembling and the details of dataset filtering and curation. Being already used by such industry leaders as Anthropic and multiple LLM training projects, we believe that Common Corpus will become a critical infrastructure for open science research in LLMs.
@@ 1 Introduction
Large Language Models have been defined by large training data. While there are several candidates for the first modern language model based on transformer architecture, including GPT-1 (Radford et al., ), ULMFIT (Howard and Ruder, 2018), or Sentence Neuron (Radford et al., 2017), it is commonly acknowledged that “large” models start with GPT-3 (Brown et al., 2020). Requiring a corpus of 300 billion tokens, GPT-3 introduced a standard training data pipeline shared by nearly all language models to date: large-scale processing of web datasets (45 TB of compressed source data from Common Crawl) and additional digitized sources (e.g. Books3). Until 2025, LLM training data has grown on a logarithmic curve. The latest generation of publicly documented language models including DeepSeek v3 (DeepSeek-AI, 2025), Gemma 3 (Google, 2025), Llama 4 (Meta AI, 2025) or Qwen 3 (Qwen, 2025) have been trained on 14-36 trillion tokens. Even the gaining popularity sub-category of small language models (Wang et al., 2025) relies on large amounts of training data to fit scaling laws: Qwen 3 0.6B was trained on 36 trillion tokens, about 3,000 times more data than recommended by the original Chinchilla laws (Hoffmann et al., 2022).
As data curation also became a primary concern, the collection, maintenance, processing, and filtering of data became one of the primary cost of standard language model training. This does not factor in even larger hidden costs: the negative externalities affecting competing markets, the digital commons, and society at large.
@@
While data scraped from the web is publicly available, it is not always in the public domain. Most web data do not have sufficient metadata to determine whether it is permissively licensed. NLP practitioners have relied on the protection of fair use, claiming that the transformative nature of the use of the data allows them to use this data to train language models. There are increasingly more legal challenges to the use of this data. The New York Times sued Open AI for copyright infringement, alleging that Open AI trained their models on their articles (Roth, 2023, Pope, 2024). Due to concerns about indirect commercial exploitation, many rightholders have implemented either hard technical measures or legal provisions against model training. In 2024, it was estimated that for Terms of Service crawling restrictions, a full 45% of C4 is now restricted (Longpre et al., 2024b) and 5% is fully blocked for scraping with a disproportionate impact over quality sources (Longpre et al., 2024b). Restrictions not only affect language model pre-training but also the quality of search engine indexation and a variety of research projects analyzing and collecting content at scale. Even projects dedicated to knowledge access have faced significant pressure from AI crawlers and implemented protections that negatively impact access and user experience.
@@
These legal uncertainties have significantly impeded the development of open science research on LLMs. Previously reproducible research artifacts have been removed or taken down, impacting pre-training data, continuous pre-trained models, and evaluation datasets. Books3, which has been used in datasets like the Pile (Gao et al., 2020) and by multiple AI labs, including Meta and Microsoft, faced legal challenges (Brittain, 2023), and the original dataset was ultimately removed (Van der Sar, 2023). The LAION dataset was demonstrated to contain CSAM (Birhane et al., 2021, Thiel, 2023), and then was taken down (LAION, 2023). Later, it was re-released once suspected CSAM was removed (LAION, 2024). The Dutch-focused model GEITje was taken down (Rijgersberg, 2025) due to complaints about it having been trained on the Dutch Gigacorpus, in order to avoid legal disputes. Finally, the widely used benchmark, the Mathematics Aptitude Test of Heuristics (MATH) dataset (Hendrycks et al., 2021), was removed from Hugging Face via a DMCA takedown. All of these artifacts, which were released to further open development and evaluation of language models, were removed suddenly, making previous work unreplicable. These takedowns and legal challenges also represent a sizeable loss of investment for developers, who are often independent or small research organizations.
@@
In part, as a reaction to the use of publicly available but not permissively licensed data, web text is becoming increasingly more difficult to acquire and use. In an analysis of popular datasets such as C4 (Raffel et al., 2020), RefinedWeb (Penedo et al., 2023), and Dolma (Soldaini et al., 2024), Longpre et al. (2024c) found that just in the last year, 5% of all tokens in C4 now have restricted use, with a disproportionate number of those tokens coming from the best-maintained, most critical sources. This is largely due to changes in content owners’ and hosts’ preferences, which are changing to no longer allow scraping, especially for the purposes of training AI models.
@@
Since 2024, several initiatives have emerged to collect open data with clear licensing and in the English language. This includes: C4C, Open License Corpus, a 228 billion token corpus from a mix of public domain texts and open source code under free licenses not requiring attribution (Min et al., 2024), KL3M, a 1.2 trillion tokens corpus of administrative texts and structured data mostly from the US federal public domain (Bommarito et al., 2025), Common Pile, an ongoing data collection of 1 trillion tokens from a variety of recent sources, including a filtered version of Common Crawl data (Creative Commons Common Crawl) (Kandpal and Raffel, 2025). These projects are all monolingual, restricting in effect the reach of open-data language models to the English-speaking audience. In contrast, the most ambitious multilingual collection of permissive content pre-dates Large Language Models: C4C (2016), containing 12 million web pages in more than 50 languages filtered by Creative Commons Licenses (Habernal et al., 2016).
@@
In this paper, we present Common Corpus, which has been an ongoing effort since 2024 (see Appendix A for the development history). It has grown to become the largest open dataset for pre-training Large Language Models at about 2 trillion tokens and the only one in its size range available in multiple languages. Thus, Common Corpus is a significant contribution to the AI Commons. Through the release of this dataset, we show that open LLM research and development is possible while meeting legal and regulatory requirements—even in the EU, where AI regulations are currently the strictest. Here, we detail the composition of Common Corpus (Section 3), as well as the data provenance, license verification, and further processing (Section 4). Next, we describe the early impact of Common Corpus on the AI community (Section 5) and highlight the limitations of the work that has been done at the time of publication. Despite its size, Common Corpus is still far from covering the entire range of available resources for pre-training commons: we attribute this discrepancy to an open data paradox as major sources of open content are paradoxically little visible online and even more so in the leading sources currently used for pre-training. By describing the unique challenges coming with the aggregation of large open source, we aim to inspire further responsible data collection initiatives.
Figure 1: A schematic world map of languages in Common Corpus with a log-scaled distribution of document counts. For each language, we chose a city that is located in the region where this language is most specific to. To avoid outliers, we show only languages with 1000+ documents.
@@ 2 About Common Corpus
When talking about Common Corpus data, we use the word “open” in the strongest possible sense. Not only is the data available, but we also provide essential details about the data provenance, data processing, and important information about the contents of each dataset. Following the Open Source Initiative’s definition of open-source AI, our data is open in terms of openness of use, meaning that use is permitted for “any purpose and without having to ask for permission” (Open Source Initiative, 2024). To achieve this, models must be trained on datasets that are free from copyright or other relevant legal limitations. This is currently a limitation of existing open datasets for training LLMs.
Common Corpus, therefore, provides valuable training tokens that will not be subject to the same restrictions. Additionally, the data in Common Corpus are different from other corpora, primarily composed of web text. Common Corpus contains multilingual data in a variety of high- and low-resource languages (see Figure 1 for language distribution), covering diverse genres, time periods, and domains (in Section 3, we detail each part of the dataset). Therefore, Common Corpus contributes to data diversity in the open pre-training data ecosystem. This is important for developing powerful and generalizable model performance. Common Corpus can be used on its own or in conjunction with existing open datasets, according to one’s needs and the desired use case of a language model.
@@
(b) A two-component t-SNE visualization of a subset of the Common Corpus.

### v2

@@ Abstract
Large Language Models (LLMs) are pre-trained on large data from different sources and domains. These datasets often contain trillions of tokens, including large portions of copyrighted or proprietary content, which raises questions about the legal use of such models. This underscores the need for truly open pre-training data that complies with data security regulations. In this paper, we introduce Common Corpus, the largest open dataset for LLM pre-training. The data assembled in Common Corpus are either uncopyrighted or under open licenses and amount to about two trillion tokens. The dataset contains a wide variety of languages, ranging from the high-resource European languages to some low-resource languages rarely represented in pre-training datasets. In addition, it includes a large amount of code data. The diversity of data sources in terms of covered domains and time periods opens up the paths for both research and entrepreneurial needs in diverse areas of knowledge. In this paper, we present the detailed provenance of data assembling and the details of dataset filtering and curation. We train two small language models on Common Corpus and find that they perform comparably to other models of their size, indicating that our dataset is suitable for multilingual pretraining. Common Corpus represents a key contribution to the ecosystem for open science research on Large Language Models.
[Uncaptioned image] PleIAs/common_corpus
@@ 1 Introduction
Large Language Models demand large amounts of training data. GPT-3 (Brown et al., 2020) is considered the first “large” language model. Trained on 300 billion tokens, GPT-3 introduced a standard training data pipeline shared by nearly all language models to date: large-scale processing of web datasets (45 TB of compressed source data from Common Crawl) and additional digitized sources (Books3). Until 2025, LLM training data sizes have grown logarithmically. The latest generation of publicly documented language models including DeepSeek v3 (Liu et al., 2025), Gemma 3 (Kamath et al., 2025), Llama 4 (Meta, 2025) or Qwen 3 (Yang et al., 2025) have been trained on 14-36 trillion tokens. Even “small language models” (Wang et al., 2025) rely on large amounts of training data to fit scaling laws: Qwen 3 0.6B was trained on 36 trillion tokens, which is 3,000 times more than Chinchilla-optimal estimates (Hoffmann et al., 2022).
Figure 1: Proportional treemap of Common Corpus collections and their most popular languages.
@@
As data quality has increasingly been an area of concern, the collection, maintenance, processing, and filtering of data has become one of the main costs in language model training. But data curation at scale may also have hidden costs: negative externalities affecting competing markets, the digital commons, and society at large. While data scraped from the web is publicly available, it is not always in the public domain. Most web data does not have sufficient metadata to determine whether it is permissively licensed. NLP practitioners have relied on the protection of fair use, claiming that the transformative nature of the use of the data allows them to use this data to train language models. There are increasingly more legal challenges to the use of this data. The New York Times sued OpenAI for copyright infringement, alleging that OpenAI trained their models on NYT articles  (Roth, 2023; Pope, 2024). Due to concerns about indirect commercial exploitation, many rightholders have implemented either hard technical measures or legal provisions against model training. In 2024, it was estimated that for Terms of Service crawling restrictions, a full 45% of C4 is now restricted (Longpre et al., 2024b) and 5% is fully blocked for scraping with a disproportionate impact over quality sources (Longpre et al., 2024b). Restrictions not only affect LLM pre-training but also the quality of search engine indexation and a variety of research projects analyzing and collecting content at scale. Even projects dedicated to knowledge access have faced significant pressure from AI crawlers and implemented protections that negatively impact access and user experience.
@@
Legal uncertainties have significantly impeded the development of open science research on LLMs. Previously reproducible research artifacts have been removed or taken down, impacting pre-training data, continuous pre-trained models, and evaluation datasets. Books3, which has been used to train numerous models, faced legal challenges (Brittain, 2023). The original dataset was ultimately removed due to a DMCA take-down (Van der Sar, 2023). The LAION dataset was demonstrated to contain CSAM (Birhane et al., 2021; Thiel, 2023), and taken down (LAION, 2023), and then re-released once suspected CSAM was removed (LAION, 2024). The Dutch model GEITje was taken down (Rijgersberg, 2025), due to complaints about training on the Dutch Gigacorpus, in order to avoid legal disputes. Finally, the widely used benchmark, the Mathematics Aptitude Test of Heuristics (MATH) dataset (Hendrycks et al., 2021), was removed from Hugging Face via a DMCA takedown. All of these artifacts, which were released to further open development and evaluation of language models, were removed suddenly, making previous work unreplicable. These takedowns and legal challenges also represent a sizeable loss of investment for developers, who are often independent or small research organizations.
@@
In part as a reaction to the use of publicly available but not permissively licensed data, web text is also becoming harder to acquire and use. In an analysis of popular datasets such as C4 (Raffel et al., 2020), RefinedWeb (Penedo et al., 2023), and Dolma (Soldaini et al., 2024), Longpre et al. (2024b) found that just in the last year, 5% of all tokens in C4 now have restricted use, with a disproportionate number of those tokens coming from the best-maintained, most critical sources. This is largely due to changes in content owners’ and hosts’ preferences, which are changing to no longer allow scraping, especially for the purposes of training AI models.
@@
Since 2024, several initiatives have emerged to collect open data in English with clear licensing. This includes: C4C (Habernal et al., 2016); Open License Corpus (Min et al., 2024), a 228 billion token corpus from a mix of public domain texts and open source code under free licenses; KL3M (Bommarito et al., 2025), a 1.2 trillion tokens corpus of administrative texts and structured data mostly from the US federal public domain; Common Pile (Kandpal et al., 2025), a data collection of 1 trillion tokens from a variety of recent sources, including a filtered common crawl (Creative Commons Common Crawl). All these projects are monolingual, restricting in effect the reach of language models to the English-speaking audience. In contrast, the most ambitious multilingual collection of permissive content predates Large Language Models: C4C contains 12 million web pages in more than 50 languages filtered by Creative Commons Licenses (Habernal et al., 2016).
@@
Common Corpus has grown to become the largest fully open pre-training dataset at about 2 trillion tokens and the only one in its size range having high multilingual diversity. Through this release, we show that open LLM research and development is possible while meeting legal and regulatory requirements — in compliance with even the strictest AI regulations, such as in the European Union. In this paper, we detail the composition of Common Corpus and the process of data collection and curation, and license clearing. Despite its size, Common Corpus is still far from covering the entire range of available resources. We attribute this discrepancy to an open data paradox as major sources of open content are paradoxically little visible online and even more so in the leading pre-training sources. By describing the unique challenges coming with the aggregation of large open source, we aim to inspire further initiatives. We also train two small language models on our dataset and find that it offers comparable performance to existing multilingual models.
@@ 2 About Common Corpus
When talking about Common Corpus data, we use the word “open” in the strongest possible sense. Not only is the data available, but we also provide essential details about the data provenance, data processing, and important information about the contents of each dataset. The Open Source Initiative has also defined open-source AI in terms of openness of use, where open means that use is permitted for “any purpose and without having to ask for permission” (Open Source Initiative, 2024). To achieve this, models must be trained on datasets that are free from copyright or other legal limitations. This is currently a limitation of existing open datasets for training LLMs.
Figure 2: A schematic world map of languages in Common Corpus with a log-scaled distribution of document counts. For each language, we chose a city that is located in the region where this language is most specific to. To avoid outliers, we show only languages with 10,000+ documents.
Common Corpus, therefore, provides valuable training tokens that will not be subject to the same restrictions. Additionally, the data in Common Corpus are different from other corpora, primarily composed of web text. It contains multilingual data in a variety of high- and low-resource languages (see Figure 2 for language distribution), covering diverse genres, time periods, and domains (in Section 4, we detail each part of the dataset). Therefore, it contributes to data diversity in the open pre-training data ecosystem. This is important for developing powerful and generalizable model performance. Common Corpus can be used on its own or in conjunction with existing open datasets, according to one’s needs and the desired use case of a language model.
Common Corpus was developed with consideration for developing best practices for open-source LLM development (The AI Alliance, 2024; Longpre et al., 2024a; Duprieu and Berkouk, 2024; Baack et al., 2025). We highlight our adherence to those suggested by Baack et al. (2025):
  • Provide useful documentation. We provide information about dataset provenance and processing (Sections 4 and 5) and share key statistics to help potential users understand the applications of the dataset. Dataset documentation improves reproducibility, helps prevent misuse, and aids downstream users to best utilize the dataset (Longpre et al., 2024a).
  • Do not capture highly sensitive data. We remove personally identifiable information from our datasets (Section 4).
Common Corpus is composed of six collections: Open Government, Open Culture, Open Science, Open Web, Open Code, and Open Semantic (Figure 1), and in total amounts to 1,998,647,168,282 tokens [^1] The token counts are calculated using PleIAs tokenizer trained on a representative subsample of Common Corpus, available at: [Uncaptioned image] PleIAs/Pleias-350m-Preview. The token counts in each collection are listed in Appendix B. To highlight the diversity, we visualize the timeline of the collected documents and embeddings of a subsample in Figure 3. Each collection comprises multiple datasets, for which we provide details on provenance and other key information in the corresponding subsections. Each data object contains a license, language(s), a collection or domain of specialization, and other metadata, allowing one to filter out a desired subset. We note that the majority of the data in Common Corpus is in the public domain (see Appendix B).
@@
(b) A two-component t-SNE visualization of subsets of Common Corpus collections, C4, and FineWeb.